OpenStudy (anonymous):

How can I find the derivative of the following function? f(x) = 5arctan(xe^x)

jimthompson5910 (jim_thompson5910):

if y = arctan(x), then dy/dx = 1/(1+x^2)

jimthompson5910 (jim_thompson5910):

you will use the chain rule first, then the product rule will be used later on

jimthompson5910 (jim_thompson5910):

let me know if that helps or not to get started

OpenStudy (anonymous):

@jim_thompson5910 how do you get 1/(1+x^2) ? Is that just something i have to memorize? For example like how the derivative of sinx = cosx ?

OpenStudy (rational):

you may memorize but derivation isn't that hard

OpenStudy (rational):

\[\arctan x = y \implies x = \tan y \implies \frac{dx}{dy}=\sec^2y \] \[\implies \frac{dy}{dx}=\cos^2y = \frac{1}{1+x^2}\]

OpenStudy (anonymous):

Oh so i can use a triangle to find the derivative, correct? Thats what you did above right?

OpenStudy (rational):

Exactly!

OpenStudy (anonymous):

Okay I understand! I'm gonna try to derive this given what I understood, and I'll tell you my answer :D i feel like its gonna be wrong though xD

OpenStudy (rational):

ok sure good luck!

OpenStudy (anonymous):

Ok so this is what I did... \[\tan(y) = xe^x\] \[\frac{ dy }{ dx }\tan(y) = \frac{ dy }{ dx } xe^x\] \[\sec^2 yy' = e^x +xe^x\] \[y'=\frac{ e^x + xe^x }{ \sec^2 y } = e^x \frac{ 1+x }{ \sqrt{1+xe ^{2x}} }\] \[f'(x) = 5e^x \frac{ 1+x }{ \sqrt{1+xe ^{2x}} }\]

OpenStudy (anonymous):

Is that right? I can simplify it more right?

OpenStudy (anonymous):

I can't think of a way to simplify the denominator

OpenStudy (amilapsn):

please check what you've done for \(sec^2y\)...

OpenStudy (anonymous):

Ohhh I forgot to square it!! Right??

OpenStudy (amilapsn):

hm

OpenStudy (anonymous):

So then it would be \[y'=5 \frac{ e^x(1+x) }{ 1+2xe ^{2x}+x^2e ^{4x} }\]

OpenStudy (amilapsn):

no

OpenStudy (anonymous):

Oh no no no wait

OpenStudy (anonymous):

\[y'=5e^x \frac{ 1+x }{ 1+xe ^{2x} }\]

OpenStudy (amilapsn):

:o)

OpenStudy (anonymous):

Hahaha right?? XD

OpenStudy (anonymous):

But then what can I do next? Can I simplify it any more?

OpenStudy (anonymous):

Nope I'm still wrong!

OpenStudy (amilapsn):

hm

OpenStudy (amilapsn):

i didn't notice it..

OpenStudy (anonymous):

The part when I used the Pythagorean theorem to get \[\sqrt{1+xe ^{2x}}\] was wrong! It should be \[\sqrt{1+x^2e ^{2x}}\]

OpenStudy (amilapsn):

hm

OpenStudy (anonymous):

So at the end I get \[y'=5e^x \frac{ 1+x }{ 1+x^2e ^{2x} }\]

OpenStudy (anonymous):

Cuz \[(xe^x)^2 = x^2e ^{2x}\] and not \[xe ^{2x} \] is that right?

OpenStudy (amilapsn):

right.

OpenStudy (anonymous):

Awesome!!! Phewww haha that was such a struggle!! Thanks so so much @amilapsn and @rational

OpenStudy (amilapsn):

u're welcome.